OpenStudy (anonymous):

et f(x) = 5 x^2 + 4 x - 8. Answer the following questions. 1. Find the average slope of the function f on the interval [-1,1]. Average Slope:= ___________ 2. Verify the Mean Value Theorem by finding a number c in (-1,1) such that f'(c) = __________ Answer: c =

OpenStudy (campbell_st):

so once you have th answer to part 1 above you need f'(x) = 10x + 4 so f'(c) = 10c + 4 applying the mean value theorem you have \[f'(c) = \frac{f(1) - f(-1)}{1 - (-1)} \] which means \[10c + 4 = \frac{f(1) - f(-1)}{1 - (-1)}\] now just solve for c

OpenStudy (campbell_st):

and for part 1 you should get f(-1) = 5 - 4 - 8 = -7 f(1) = 5 + 4 - 8 = 1 so using \[avg..slope = \frac{f(1) - f(-1)}{1 - (-1)}\] \[Avg.. slope = \frac{1 --7}{1 - -1}\] then you find the average slope is Avg Slope = 4

OpenStudy (anonymous):

Thanks for the help it makes more sense now :)

OpenStudy (campbell_st):

sorry that I did it back the front... but I'm glad it makes sense